OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), -2x+7y=-1 4x-3y=-19

OpenStudy (mathstudent55):

Multiply the first equation by 2 and add it to the second equation. The x variable will cancel out and you can solve for y. Then substitute the value of y into the first original equation, and solve for x. Then write the answer as (x, y) using the values you got for x and y.

OpenStudy (anonymous):

for y i got -11/21 and x i got -61/14 and that sounds wrong

OpenStudy (mathstudent55):

-2x+7y=-1 4x-3y=-19 Multiply 1st eq by 2 and write second eq below it: -4x + 14y = -2 4x - 3y = -19 Add the equations: 11y = -21 Divide both sides by 11: y = -21/11 How did you get y = -11/21?

OpenStudy (anonymous):

i must have mixed the two together by accident thank you

OpenStudy (anonymous):

so as an answer i would get (-68/11,-21/11)

OpenStudy (mathstudent55):

I'll continue: Now that we have y = -21/11, substitute it into first original equation: -2x+7y=-1 -2x + 7(-21/11) = -1 -2x - 147/11 = -1 -2x = -1 + 147/11 -2x = -11/11 + 147/11 -2x = 136/11 x = -68/11 Solution: (-68/11, -21/11)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (anonymous):

ok cool thank you i was just making sure